Find the value of x and y if: 5x -3y = 35 and y = x-y
Answer by checkley75(3666)

5x -3y = 35
y = x-y or x=2y
5(2y)-3y=35
10y-3y=35
7y=35
y=35'7
y=5 ans.
x=2*5
x=10 ans.
Proof
5*10-3*5=35
50-15=35
35=35